The tax-inclusive price of a carton of cigarettes in Canada is $100 per carton. Per capita consumption is 10 cartons. Federal and provincial taxes are currently $50 per carton.

Suppose the two levels of government wish to increase the cigarette tax by 20 percent. If the price elasticity of demand is -0.25, what is the additional per capita tax revenue that the two levels of government can expect to collect?
Answer #1

In the initial situation

Per capita consumption=Q1=10 cartons

Price per carton=P1=$100

Tax per carton=T1=$50

Basic Price per carton=B=P1-T1=100-50=$50

Per capita tax revenue=TR1=Q1*T1=10*50=$500

In the revised situation

Tax per carton=T2=$50*(1+20%)=$60

Price per carton=P2=B+T2=50+60=$110

Percent change in price=(110-100)/105=10%

Percent change in quantity demanded=Price elasticity of demand*Change in price=-0.25*10%=-2.5%

Per capita consumption=Q2=Q1*(1+Change in quantity demanded)= 10*(1-2.5%)=9.75 cartons

Per capita tax revenue=TR2=Q2*T2=9.75*60=$585

Additional per capita tax revenue=TR2-TR1=585-500=$85